# Predictive Coding Networks and Inference Learning: Tutorial and Survey

### v1

@@ 4.2. IL in its Natural Regime
Although most works study properties of the learning algorithm as a whole, the work of (frieder_non-convergence_2022, ) showed a useful result concerning the inference phase. Using a dynamical systems perspective, they formally prove that the inference phase (during training and testing) converges only when $\gamma$ and $\alpha$ (the inference and learning rate, respectively) are smaller than 1. This is always done in practice, but (frieder_non-convergence_2022, ) gave theoretical assurance that this is indeed the case. Then, the first contribution to a theory for why IL sometimes performs better than BP was provided by (song_inferring_2024, ), which according to the authors lies in reduced weight interference. ‘Catastrophic interference’ is a well-known pathology of BP to abruptly and drastically forget previously learned information upon learning new information (kirkpatrick_overcoming_2017, ). According to (song_inferring_2024, ), this is reduced in IL, since before weight update, activation nodes are changed to a new state which is closer to what the network should ideally predict to obtain the observed outcome. They call this mechanism prospective configuration, since IL is ‘prospective’ in the sense that it can foresee side effects of weight modifications, leading to decreased weight interference. This helps to explain improved empirical results seen in continual learning tasks (which are a challenge precisely due to this interference), online learning, and the generally observed faster convergence.
@@
In a parallel line of work, (alonso_theoretical_2022, ) showed that IL approximates implicit SGD, which provides increased stability across learning rates. This is empirically what (alonso_theoretical_2022, ) observe for IL. The authors also develop a variant of PC in which updates become equal to implicit SGD, further improving the stability. This result is extended in (alonso_understanding_2024, ), where the authors observe that implicit SGD is sensitive to second-order information, i.e. the second derivative of the loss landscape (its curvature). This can speed up convergence. As such, when IL approximates implicit SGD, IL uses this information, which explains the faster convergence also observed by (song_inferring_2024, ).
@@
This connects to (innocenti_understanding_2023, ), which related IL to trust region (TR) methods. These are a well-known method in optimization, which can be contrasted with line-search methods such as gradient descent (yuan_recent_2015, ). Such methods define a neighborhood around the current best solution as a trust region in each step, which is changed adaptively in addition to the step direction (yuan_recent_2015, ). The authors show (innocenti_understanding_2023, ) that IL’s inference phase of PC can be understood as solving a TR problem on the BP loss, using a trust region defined by a second derivative of the energy. By using this second-order information, trust region methods are well-known to be better at escaping saddle points. Indeed, (innocenti_understanding_2023, ) show that these properties transfer to IL, thus drawing a similar conclusion as (alonso_understanding_2024, ). They prove on toy problems that IL escapes saddle points faster, and provide evidence that this happens in larger networks by training deep chains, where IL converges significantly faster than BP. These results help to explain the empirical results showing faster convergence of PC compared to BP (alonso_theoretical_2022, ; song_inferring_2024, ).
Theoretical results of (millidge_theoretical_2022, ) are concerned with the convergence in PCN’s energy landscape. The authors showed that that when activity nodes are initialized with the so-called ‘energy gradient bound’, IL converges to the same extrema in the loss as BP. This can be understood by writing PCN’s energy as a sum of two terms $E=\mathcal{L}+\widetilde{E}$, where $\mathcal{L}=(\bm{\epsilon}^{L})^{2}$ is the energy of the last layer, equal to the BP MSE loss, and $\widetilde{E}=\sum_{\ell=1}^{L-1}(\bm{\epsilon}^{\ell})^{2}$ is the residual energy. Now, during inference, minimizing $E$ is guaranteed to minimize $\mathcal{L}$ if the following holds (cf. (millidge_theoretical_2022, ) for a derivation):
(37) $-\left(\partialderivative{\mathcal{L}}{\bm{a}^{\ell}}\right)^{\!\top}\partialderivative{\widetilde{E}}{\bm{a}^{\ell}}\leq\left\|\partialderivative{\mathcal{L}}{\bm{a}^{\ell}}\right\|^{2}~{},$
@@
which implies that the gradient of $\mathcal{L}$ is greater than the negative gradient of the residual energy. At the beginning of inference this always holds when using feedforward-pass initialization of hidden activities, since then ${\partial\widetilde{E}}/{\partial\bm{a}^{\ell}}=0$. Indeed, (millidge_theoretical_2022, ) show that this keeps being true also in practice throughout inference, with a sufficiently small learning rate (up to 0.1, which is 2-3 orders of magnitude larger than step sizes typically used in practice). If the learning step is additionally written as a minimization of $\mathcal{L}$, (millidge_theoretical_2022, ) show that IL will lead to the same minima as BP, even though the weight updates of BP and IL are different. This helps explain the fact that in practice, performance of IL very closely matches BP (whittington_approximation_2017, ; alonso_theoretical_2022, ; rosenbaum_relationship_2022, ; salvatori_stable_2023, ; song_inferring_2024, ). Consequently, deep PCNs could potentially achieve the same generalization performance as BP while maintaining their advantages.
@@
This result seems to be contradicted in part by (frieder_bad_2024, ) however, which shows that use of activation functions with many zeros (such as ReLU) can be problematic for PCNs, since they can prevent convergence of learning. This result is interesting in light of (innocenti_understanding_2023, ) which highlights how ReLU does not show the same speed benefits as other activation functions. Interestingly, even for ANNs, more sophisticated treatments indicate that ReLU is a suboptimal choice of activation function due to the fact that the critical regime reduces to a single point in phase space (roberts_principles_2022, ). In general, these works show how although PCNs and ANNs are similar in many respects, one cannot rely completely on existing knowledge about BP-based frameworks for predicting properties of PCNs.
@@ 4.3. Improving Performance
As was mentioned, the notion of doing EM incrementally was a recent successful approach to improving the efficiency of IL, improving the time complexity of a weight update by a factor $T$. A second proposal for modifying IL to improve its complexity comes from (alonso_understanding_2024, ), who suggest that the inference phase includes a delay in the computation of errors, where updates in earlier layers are zero at the beginning of inference. They suggest changing the inference to not update layers in parallel, but sequentially from layer $L$ to layer 00, thus proposing ‘sequential inference’. The authors observe that this decreases the number of time-steps required per weight update without sacrificing performance. However, we remark that this change effectively updates lower layers faster than would otherwise happen, and is thus a change to (11). It is also reminiscent of Z-IL as discussed above, making IL closer to BP in some sense. As such, it is unclear whether this preserves the advantages of IL observed in other works. In the same work, the authors build on the observation of (alonso_theoretical_2022, ) that IL sometimes converges to poor local minima when trained with SGD. They suggest this may be caused by increasingly small weight updates in lower layers. They suggest an optimizer created for IL specifically (dubbed Matrix Update Equalization (MQ)-optimizer) that requires less memory than Adam, and similarly appears to avoid poor local minima, with promising initial results.
@@ 5. Conclusion
Despite the remarkable successes of deep learning witnessed in recent years, biological learning is still superior to machine learning (ML) in many key respects, such as energy and data efficiency (zador_catalyzing_2023, ). Whereas the most successful deep learning approaches rely on backpropagation (BP) as a training algorithm, recent work (song_inferring_2024, ) has identified learning mechanisms in predictive coding networks (PCNs), trained with inference learning (IL), to provide an improved model for many features of biological learning.
Research on PCNs has recently witnessed a burst in activity. PCNs build on the neuroscientific theory of Predictive Coding (PC), thus exemplifying recent calls for renewed emphasis on neuroscience-inspired methods to artifical intelligence, called NeuroAI (zador_catalyzing_2023, ). PCNs are trained with Inference Learning and entail a flexible framework for ML that goes beyond traditional (BP-trained) Artificial Neural Networks (ANNs). Despite the recent burst in activity, a comprehensive and mathematically oriented introduction aimed at machine learning practitioners has been lacking, which this tutorial has attempted to provided.
This work has discussed three complementary perspectives on PCNs (see fig. 3):
@@
    a model that generalizes the structure of ANNs (the most recent notion, due to (salvatori_learning_2022, ));
@@
    a probabilistic latent variable model for unsupervised learning (its inception (rao_predictive_1999, ));
@@
    a learning algorithm (IL) that can be compared to BP (the most common conception in the recent literature, as for example in (song_inferring_2024, )).
@@
The first perspective, PCNs as a generalized ANN, is justified by the fact that mathematically, the structure of PCNs forms a superset of traditional ANNs (a fact that follows from earlier work, but has not been pointed out as a general conception of PCNs). Thus, we expect PCNs share the appealing property of being universal function approximators, while also providing a broader, more general framework as a neural network. Namely, a simple choice of direction gives networks appropriate not only for supervised learning and classification, but also unsupervised learning in generative models. Next to this, the generalization of PCNs to arbitrary graphs (PC graphs) means a large new set of structures untrainable by BP can be studied theoretically and empirically. Given the topological nature of several advances in machine learning in the past (such as residual networks and the notion of depth itself), this suggests an interesting avenue for further work.
@@
The second perspective, PCN as probabilistic latent variable model, provides a principled derivation of equations that underlie PCNs. In addition, it brings to light the mentioned connections to other ML algorithms. PCNs share several properties both with classic methods such as factor analysis, but also to modern methods such as diffusion models – associated with the remarkable success of generative AI witnessed in recent years. In general, PCNs’ probabilistic nature and these various connections also suggest they could potentiallly be developed further as a method for Bayesian deep learning, which aim to reliably assess uncertainties and incorporate existing knowledge (papamarkou_position_2024, ).
@@
The third perspective has been the focus of most of the recent literature on PCNs. Existing implementations of IL are still inefficient compared to BP, which has prevented adoption of PCNs for practical applications, but recent work has taken important steps towards improving efficiency, showing that IL no longer scales with the depth of the network (i.e. it is an order $L$ network depth faster per weight update), if sufficient parallelization is achieved. The reason for this is IL’s local computations, which could bring advantages as deep networks keep scaling up, or on neuromorphic hardware. Barring the issue of efficiency, IL demonstrates favorable properties which call for further investigation. In particular, decreased weight interference and sensitivity to second-order information, which can lead to faster convergence and increased performance on certain learning tasks (e.g. continual learning and adaptation to changing environments). To what extent these results generalize and scale remains to be studied, but theoretical results suggest that IL should perform at least as well as BP in terms of generalization, potentially opening the door to additional advantages.
@@
In summary, this tutorial provides a formal overview of PCNs that can accelerate future work by the ML community. Complementary to existing surveys and reviews, it has provided a mathematically oriented introduction, using detailed formal specification and connections to various other methods in ML, while remaining accessible, by clarifying and structuring different perspectives on PCNs. As such, it can be a useful starting point for researchers interested in PCNs, and can help them pursue further work on this fascinating model in the emerging field of NeuroAI.
@@ A.1. Expectation Maximization
where we defined $\operatorname{NLL}(\theta)$ as the negative log-likelihood cf. (17). Many deep learning models increase their expressivity through latent (hidden) variables $\bm{z}$, low-dimensional compared to $\bm{x}$. The idea of such variables is that they capture information that is not directly observable in $\bm{x}$, and they have a simpler probability distribution by design. Latent variable models underlie much of the power of modern deep learning methods, in particular generative models (goodfellow_deep_2016, ; prince_understanding_2023, ).
@@
This integral is often intractable (prince_understanding_2023, ). Indeed, for PC this is also the case, which can be observed by filling in $p_{\theta}(\bm{x},\bm{z})={\exp}(-E_{\theta}(\bm{x},\bm{z}))$, and with the energy of Rao & Ballard, cf. (27) one has
@@
which is generally intractable due to the nonlinearity in the integrand. The same holds for the multi-layer network with (33), where the integral becomes even more complicated. Hence, we cannot just minimize $\text{NLL}(\theta)$ directly. For general generative models, this can be solved in different ways – commonly through MCMC methods, or variational inference (goodfellow_deep_2016, ). Expectation Maximization (EM), presented in (A.9) and (A.10) can be seen as a special case of variational inference (dempster_maximum_1977, ; bishop_pattern_2006, ). Here, we shall motivate EM in detail, highlighting the different steps required to get to IL.
@@
First, introduce some distribution $q_{\phi}(\bm{z})$ over the latent variables. Given a datapoint $\bm{x}^{(n)}$, EM proceeds in two steps (mehta_high-bias_2019, ; dempster_maximum_1977, ). (1) E-step: given $\bm{x}$ and the current parameter $\theta$, construct a new distribution $q_{\phi}(\bm{z})$ of latent variables, using the current posterior $p_{\theta}(\bm{z}|\bm{x})$. (2) M-step: re-estimate the parameter $\theta$ to be those with maximum likelihood, assuming $q_{\phi}(\bm{z})$ found in the previous step is the true distribution of $\bm{z}$. In other words:
(A.4) $q_{\phi}(\bm{z})$ $=p_{\theta}(\bm{z}|\bm{x})$ (E-step)
@@
Taking the $k$-means clustering algorithm as an example, these two steps correspond to assigning points to clusters, and defining clusters based on points, respectively (mackay_information_2003, ). This process works because one can show that each EM iteration decreases the true NLL, or at worst leaves it unchanged (dempster_maximum_1977, ). This can be shown by writing both steps as minimization of a single functional: the variational free energy.
@@
$=D_{\text{KL}}[q_{\phi}(\bm{z})||p_{\theta}(\bm{z}|\bm{x})]-\ln p_{\theta}(\bm{x})~{}.$
Since KL-divergences are always positive, one sees that $\mathcal{F}\geq-\ln p_{\theta}(\bm{x})$, i.e., $\mathcal{F}$ is an upper bound on the NLL. [^13] One can equivalently write $-\mathcal{F}\leq\ln p_{\theta}(\bm{x})$, i.e. $-\mathcal{F}$ as a lower bound to $\ln p_{\theta}(\bm{x})$. The marginal log-likelihood is also called the model evidence, hence $-\mathcal{F}$ is also called the evidence lower bound, or ELBO(prince_understanding_2023, ). In other words, if minimized with respect to $\theta$, the NLL is also minimized. Similarly, it can be shown that the E-step can be written as minimization of this function (neal_view_1998, ). Thus, one can write a new version of EM as follows:
@@
Using the result in (neal_view_1998, ) that EM will minimize the NLL, we return to the first formulation, (A.4) and (A.5). At first sight, the E-step still looks problematic, since evaluating the posterior $p_{\theta}(\bm{z}|\bm{x})=p_{\theta}(\bm{x},\bm{z})/p_{\theta}(\bm{x})$ requires calculating the marginal likelihood $p_{\theta}(\bm{x})$. This is analogous to computing the partition function in statistical physics, which is generally intractable. A solution is to set $q$ to a delta function, $q_{\phi}(\bm{z})=\delta(\bm{z}-\widetilde{\bm{z}})$, in which case the posterior is simply described by its maximum value. Put differently, we simplify the evaluation of the posterior in the E-step by a maximum a posteriori (MAP) estimation of $p(\bm{z}|\bm{x})$. With a delta-distribution, the expectation of (A.5) also simplifies, and we have
(A.9) $\widetilde{\bm{z}}$ $=\underset{\bm{z}}{\operatorname{argmax}}\,p_{\theta}(\bm{x},\bm{z})$ (E-step)
@@
With this choice, the parameters of $q_{\phi}(\bm{z})$ are $\phi=\widetilde{\bm{z}}$. We discuss this in more detail below.
Thus, we see having a model for $\ln p_{\theta}(\bm{x},\bm{z})$ is sufficient for minimizing the (upper bound of) the NLL using EM. By choosing a form for $p_{\theta}(\bm{x},\bm{z})$ (or $E$) one obtains different models, including PC.
@@
Writing EM as a minimization of the VFE, it can be shown that minimizations in both steps need not be complete, but can also be partial (neal_view_1998, ). Taking steps that decreases $\mathcal{F}$ (i.e., in a gradient descent formulation) will also minimize the NLL. Algorithms with partial E and/or M steps are called generalized EM algorithms (neal_view_1998, ; bishop_pattern_2006, ). Thus, we can write our process using a partial M-step:
@@
(A.13) $\Delta\bm{z}$ $\propto-\frac{dE}{d\bm{z}}$ (partial E-step)
(A.14) $\Delta{\theta}$ $\propto-\frac{dE}{d{\theta}}$ (partial M-step)

### v2

@@ 4.2. IL in its Natural Regime
Although most works study properties of the learning algorithm as a whole, the work of (Frieder and Lukasiewicz, 2022) showed a useful result concerning the inference phase. Using a dynamical systems perspective, they formally prove that the inference phase (during training and testing) converges only when $\gamma$ and $\alpha$ (the inference and learning rate, respectively) are smaller than 1. This is always done in practice, but (Frieder and Lukasiewicz, 2022) gave theoretical assurance that this is indeed the case. Then, the first contribution to a theory for why IL sometimes performs better than BP was provided by (Song et al., 2024), which according to the authors lies in reduced weight interference. ‘Catastrophic interference’ is a well-known pathology of BP to abruptly and drastically forget previously learned information upon learning new information (Kirkpatrick et al., 2017). According to (Song et al., 2024), this is reduced in IL, since before weight update, activation nodes are changed to a new state which is closer to what the network should ideally predict to obtain the observed outcome. They call this mechanism prospective configuration, since IL is ‘prospective’ in the sense that it can foresee side effects of weight modifications, leading to decreased weight interference. This helps to explain improved empirical results seen in continual learning tasks (which are a challenge precisely due to this interference), online learning, and the generally observed faster convergence.
@@
In a parallel line of work, (Alonso et al., 2022) showed both theoretically and empirically that IL approximates implicit SGD, which provides increased stability across learning rates. The authors also develop a variant of PC in which updates become equal to implicit SGD, further improving the stability. This is extended in (Alonso et al., 2024), where the authors observe that implicit SGD is sensitive to second-order information, i.e. the second derivative of the loss landscape (its curvature). This can speed up convergence. As such, when IL approximates implicit SGD, IL uses this information, which explains the faster convergence also observed by (Song et al., 2024).
@@
This connects to (Innocenti et al., 2023), which related IL to trust region (TR) methods. These are a well-known method in optimization, which can be contrasted with line-search methods such as gradient descent (Yuan, 2015). Such methods define a neighborhood around the current best solution as a trust region in each step, which is changed adaptively in addition to the step direction (Yuan, 2015). The authors show (Innocenti et al., 2023) that IL’s inference phase of PC can be understood as solving a TR problem on the BP loss, using a trust region defined by a second derivative of the energy. By using this second-order information, trust region methods are well-known to be better at escaping saddle points. Indeed, (Innocenti et al., 2023) show that these properties transfer to IL, thus drawing a similar conclusion as (Alonso et al., 2024). They prove on toy problems that IL escapes saddle points faster, and provide evidence that this happens in larger networks by training deep chains, where IL converges significantly faster than BP. These results help to explain the empirical results showing faster convergence of PC compared to BP (cf. section 2.3.5).
@@
Theoretical results of (Millidge et al., 2023b) are concerned with the convergence in PCN’s energy landscape. The authors showed that that when activity nodes are initialized with the so-called ‘energy gradient bound’, IL converges to the same extrema in the loss as BP. This can be understood by writing PCN’s energy as a sum of two terms $E=\mathcal{L}+\widetilde{E}$, where $\mathcal{L}=(\bm{\epsilon}^{L})^{2}$ is the energy of the last layer, equal to the BP MSE loss, and $\widetilde{E}=\sum_{\ell=1}^{L-1}(\bm{\epsilon}^{\ell})^{2}$ is the residual energy. Now, during inference, minimizing $E$ is guaranteed to minimize $\mathcal{L}$ if the following holds that (cf. (Millidge et al., 2023b) for a derivation) $-\left(\partialderivative{\mathcal{L}}{\bm{a}^{\ell}}\right)^{\!\top}\partialderivative{\widetilde{E}}{\bm{a}^{\ell}}\leq\left\|\partialderivative{\mathcal{L}}{\bm{a}^{\ell}}\right\|^{2}~{},$ which implies that the gradient of $\mathcal{L}$ is greater than the negative gradient of the residual energy. At the beginning of inference this always holds when using feedforward-pass initialization of hidden activities, since then ${\partial\widetilde{E}}/{\partial\bm{a}^{\ell}}=0$. Indeed, (Millidge et al., 2023b) show that this keeps being true also in practice throughout inference, with a sufficiently small learning rate (up to 0.1, which is 2-3 orders of magnitude larger than step sizes typically used in practice). If the learning step is additionally written as a minimization of $\mathcal{L}$, (Millidge et al., 2023b) show that IL will lead to the same minima as BP, even though the weight updates of BP and IL are different. This helps explain the fact that in practice, performance of IL very closely matches BP (Whittington and Bogacz, 2017; Alonso et al., 2022; Rosenbaum, 2022; Salvatori et al., 2024b; Song et al., 2024). Consequently, deep PCNs could potentially achieve the same generalization performance as BP while maintaining their advantages.
@@
This result seems to be contradicted in part by (Frieder et al., 2024) however, which shows that use of activation functions with many zeros (such as ReLU) can be problematic for PCNs, since they can prevent convergence of learning. This result is interesting in light of (Innocenti et al., 2023) which highlights how ReLU does not show the same speed benefits as other activation functions. Interestingly, even for ANNs, more sophisticated treatments indicate that ReLU is a suboptimal choice of activation function due to the fact that the critical regime reduces to a single point in phase space (Roberts et al., 2022). In general, these works show how although PCNs and ANNs are similar in many respects, one cannot rely completely on existing knowledge about BP-based frameworks for predicting properties of PCNs.
@@ 4.3. Improving Performance
As was mentioned, the notion of doing EM incrementally was a recent successful approach to improving the efficiency of IL, improving the time complexity of a weight update by a factor $T$. A second proposal for modifying IL to improve its complexity comes from (Alonso et al., 2024), who suggest that the inference phase includes a delay in the computation of errors, where updates in earlier layers are zero at the beginning of inference. They suggest changing the inference to not update layers in parallel, but sequentially from layer $L$ to layer 00, thus proposing ‘sequential inference’. The authors observe that this decreases the number of time-steps required per weight update without sacrificing performance. However, we remark that this change effectively updates lower layers faster than would otherwise happen, and is thus a change to (11). It is also reminiscent of Z-IL as discussed above, making IL closer to BP in some sense. As such, it is unclear whether this preserves the advantages of IL observed in other works. In the same work, the authors build on the observation of (Alonso et al., 2022) that IL sometimes converges to poor local minima when trained with SGD. They suggest this may be caused by increasingly small weight updates in lower layers. They suggest an optimizer created for IL specifically (dubbed Matrix Update Equalization (MQ)-optimizer) that requires less memory than Adam, and similarly appears to avoid poor local minima, with promising initial results.
@@ 5. Conclusion
Despite the remarkable successes of deep learning witnessed in recent years, biological learning is still superior to machine learning (ML) in many key respects, such as energy and data efficiency (Zador et al., 2023). Whereas the most successful deep learning approaches rely on backpropagation (BP) as a training algorithm, recent work (Song et al., 2024) identified the learning mechanisms in predictive coding networks (PCNs), trained with inference learning (IL), to provide an improved model for many features of biological learning.
Research on PCNs has recently witnessed a burst in activity. These networks build on the neuroscientific theory of predictive coding (PC), thus exemplifying recent calls for renewed emphasis on neuroscience-inspired methods to artifical intelligence, called NeuroAI (Zador et al., 2023). PCNs entail a flexible framework for ML that goes beyond traditional (BP-trained) artificial neural networks (ANNs). Despite the recent burst in activity, a comprehensive and mathematically oriented introduction aimed at machine learning practitioners has been lacking, which this tutorial has attempted to provide.
This work has discussed three complementary perspectives on PCNs (cf. fig. 3):
@@
    a model that generalizes the structure of ANNs (the most recent notion, due to (Salvatori et al., 2022a));
@@
    a probabilistic latent variable model for unsupervised learning (its inception (Rao and Ballard, 1999));
@@
    a learning algorithm (IL) that can be compared to BP (the most common conception in the recent literature, as for example in (Song et al., 2024)).
@@
The first perspective, PCNs as a generalized ANN, is justified by the fact that mathematically, the structure of PCNs forms a superset of traditional ANNs (a fact that follows from earlier work, but has not been pointed out as a general conception of PCNs). Thus, we expect PCNs share the appealing property of being universal function approximators, while also providing a broader, more general framework as a neural network. Namely, a simple choice of direction gives networks appropriate not only for supervised learning and classification, but also unsupervised learning in generative models. Next to this, the generalization of PCNs to arbitrary graphs (PC graphs) (Salvatori et al., 2022a) means a large new set of structures untrainable by BP can be studied theoretically and empirically. Given the topological nature of several advances in machine learning in the past (such as residual networks (Zagoruyko and Komodakis, 2016) and the notion of depth itself (Prince, 2023)), this suggests an interesting avenue for further work.
@@
The second perspective, PCN as probabilistic latent variable model, provides a principled derivation of equations that underlie PCNs. In addition, it brings to light the mentioned connections to other ML algorithms. PCNs share several properties both with classic methods such as factor analysis, but also to modern methods such as diffusion models – associated with the remarkable success of generative AI witnessed in recent years (Prince, 2023). In general, PCNs’ probabilistic nature and these various connections also suggest they could potentiallly be developed further as a method for Bayesian deep learning, which aim to reliably assess uncertainties and incorporate existing knowledge (Papamarkou et al., 2024; Wang and Yeung, 2020).
@@
The third perspective has been the focus of most of the recent literature on PCNs (Whittington and Bogacz, 2017; Rosenbaum, 2022; Millidge et al., 2023a). Existing implementations of IL are still inefficient compared to BP, which has prevented adoption of PCNs for practical applications, but recent work has taken important steps towards improving efficiency, showing that IL no longer scales with the depth of the network (i.e. it is an order $L$ network depth faster per weight update), if sufficient parallelization is achieved (Salvatori et al., 2024b). The reason for this is IL’s local computations, which could bring advantages as deep networks keep scaling up, or on neuromorphic hardware. Barring the issue of efficiency, IL demonstrates favorable properties which call for further investigation. In particular, decreased weight interference (Song et al., 2024) and sensitivity to second-order information (Alonso et al., 2022; Innocenti et al., 2023), which can lead to faster convergence and increased performance on certain learning tasks (e.g. continual learning and adaptation to changing environments). To what extent these results generalize and scale remains to be studied, but theoretical results suggest that IL should perform at least as well as BP in terms of generalization, potentially opening the door to additional advantages (Millidge et al., 2023b).
@@
In sum, this tutorial offers a formal overview of PCNs, complementing existing surveys with a detailed mathematical introduction and connections to other ML methods. It clarifies and structures various perspectives, making it an accessible starting point for researchers. As such, it can aid in accelerating future work in the emerging field of NeuroAI.
Acknowledgements.
B.E.v.Z. and E.L.v.d.B. thank our partners, members from the HONDA project between the Honda Research Institute in Japan and Utrecht University in the Netherlands, for supporting this research. The funders had no role in study design, data collection and analysis, decision to publish or preparation of the manuscript. B.E.v.Z. thanks Lukas Arts and Mathijs Langezaal for helpful discussions and feedback.
@@ A.1. Expectation Maximization
where we defined $\operatorname{NLL}(\theta)$ as the negative log-likelihood cf. (17). Many deep learning models increase their expressivity through latent (hidden) variables $\bm{z}$, low-dimensional compared to $\bm{x}$. The idea of such variables is that they capture information that is not directly observable in $\bm{x}$, and they have a simpler probability distribution by design (Prince, 2023). Latent variable models underlie much of the power of modern deep learning methods, in particular generative models (Goodfellow et al., 2016; Prince, 2023).
@@
This integral is often intractable (Prince, 2023). Indeed, for PC this is also the case, which can be observed by filling in $p_{\theta}(\bm{x},\bm{z})={\exp}(-E_{\theta}(\bm{x},\bm{z}))$, and with the energy of Rao & Ballard, cf. (26) one has
@@
which is generally intractable due to the nonlinearity in the integrand. The same holds for the multi-layer network with (32), where the integral becomes even more complicated. Hence, we cannot just minimize $\text{NLL}(\theta)$ directly. For general generative models, this can be solved in different ways – commonly through MCMC methods, or variational inference (Goodfellow et al., 2016). Expectation Maximization (EM), presented in (18) and (19) can be seen as a special case of variational inference (Dempster et al., 1977; Bishop, 2006). Here, we shall motivate EM in detail, highlighting the different steps required to get to IL.
@@
First, introduce some distribution $q_{\phi}(\bm{z})$ over the latent variables. Given a datapoint $\bm{x}^{(n)}$, EM proceeds in two steps (Mehta et al., 2019; Dempster et al., 1977). (1) E-step: given $\bm{x}$ and the current parameters $\theta$, construct a new distribution $q_{\phi}(\bm{z})$ of latent variables, using the current posterior $p_{\theta}(\bm{z}|\bm{x})$. (2) M-step: re-estimate the parameter $\theta$ to be those with maximum likelihood, assuming $q_{\phi}(\bm{z})$ found in the previous step is the true distribution of $\bm{z}$. In other words:
@@
[…]
Taking the $k$-means clustering algorithm as an example, these two steps correspond to assigning points to clusters, and defining clusters based on points, respectively (MacKay, 2003). This process works because one can show that each EM iteration decreases the true NLL, or at worst leaves it unchanged (Dempster et al., 1977). This can be shown by writing both steps as minimization of a single functional: the variational free energy.
@@
$=D_{\text{KL}}[q_{\phi}(\bm{z})||p_{\theta}(\bm{z}|\bm{x})]-\ln p_{\theta}(\bm{x})~{}$
(derived below, cf. section A.2.1). Since KL-divergences are always positive, one sees that $\mathcal{F}\geq-\ln p_{\theta}(\bm{x})$, i.e., $\mathcal{F}$ is an upper bound on the NLL. [^16] One can equivalently write $-\mathcal{F}\leq\ln p_{\theta}(\bm{x})$, i.e. $-\mathcal{F}$ as a lower bound to $\ln p_{\theta}(\bm{x})$. The marginal log-likelihood is also called the model evidence, hence $-\mathcal{F}$ is also called the evidence lower bound, or ELBO(Prince, 2023). In other words, if minimized with respect to $\theta$, the NLL is also minimized. Similarly, it can be shown that the E-step can be written as minimization of this function (Neal and Hinton, 1998). Thus, one can write a new version of EM as follows:
@@
Using the result in (Neal and Hinton, 1998) that EM will minimize the NLL, we return to the first formulation, (A.4) and (A.5). At first sight, the E-step still looks problematic, since evaluating the posterior $p_{\theta}(\bm{z}|\bm{x})=p_{\theta}(\bm{x},\bm{z})/p_{\theta}(\bm{x})$ requires calculating the marginal likelihood $p_{\theta}(\bm{x})$. This is analogous to computing the partition function in statistical physics, which is generally intractable. A solution is to set $q$ to a delta function, $q_{\phi}(\bm{z})=\delta(\bm{z}-\widetilde{\bm{z}})$, in which case the posterior is simply described by its maximum value. Put differently, we simplify the evaluation of the posterior in the E-step by a maximum a posteriori (MAP) estimation of $p(\bm{z}|\bm{x})$. With a delta-distribution, the expectation of (A.5) also simplifies, and we have
@@
[…]
With this choice, the parameters of $q_{\phi}(\bm{z})$ are $\phi=\widetilde{\bm{z}}$. We discuss this in more detail below. Thus, we see having a model for $\ln p_{\theta}(\bm{x},\bm{z})$ is sufficient for minimizing the (upper bound of) the NLL using EM. By choosing a form for $p_{\theta}(\bm{x},\bm{z})$ (or $E$) one obtains different models, including PC.
@@
Writing EM as a minimization of the VFE, it can be shown that minimizations in both steps need not be complete, but can also be partial (Neal and Hinton, 1998). Taking steps that decreases $\mathcal{F}$ (i.e., in a gradient descent formulation) will also minimize the NLL. Algorithms with partial E and/or M steps are called generalized EM algorithms (Neal and Hinton, 1998; Bishop, 2006). Thus, we can write our process using a partial M-step:
@@
[…]